#16 A random sample of companies in electric utilities (I), financial services (II), and food processing...

#16

A random sample of companies in electric utilities (I), financial services (II), and food processing (III) gave the following information regarding annual profits per employee (units in thousands of dollars).

I II III
49.5 55.1 39.1
43.5 25.3 37.2
32.9 41.1 10.9
27.9 29.5 32.6
38.1 39.4 15.5
36.2 42.6
20.7

Shall we reject or not reject the claim that there is no difference in population mean annual profits per employee in each of the three types of companies? Use a 1% level of significance.

(b) Find SSTOT, SSBET, and SSW and check that SSTOT = SSBET + SSW. (Use 3 decimal places.)

SSTOT =
SSBET =
SSW =


Find d.f.BET, d.f.W, MSBET, and MSW. (Use 3 decimal places for MSBET, and MSW.)

dfBET =
dfW =
MSBET =
MSW =


Find the value of the sample F statistic. (Use 3 decimal places.)


What are the degrees of freedom?
(numerator)
(denominator)

(f) Make a summary table for your ANOVA test.

Source of
Variation
Sum of
Squares
Degrees of
Freedom
MS F
Ratio
P Value Test
Decision
Between groups NA NA.
Within groups
Total
Answer #1
I II III
count, ni = 7 5 6
mean , x̅ i = 35.543 38.08 29.65
std. dev., si = 9.581 11.592 13.225
sample variances, si^2 = 91.800 134.372 174.899
total sum 248.8 190.4 177.9 617.1 (grand sum)
grand mean , x̅̅ = Σni*x̅i/Σni =   34.28
square of deviation of sample mean from grand mean,( x̅ - x̅̅)² 1.586 14.415 21.468
TOTAL
SS(between)= SSB = Σn( x̅ - x̅̅)² = 11.105 72.073 128.807 211.9849
SS(within ) = SSW = Σ(n-1)s² = 550.797 537.488 874.495 1962.780

no. of treatment , k =   3
df between = k-1 =    2
N = Σn =   18
df within = N-k =   15
  
mean square between groups , MSB = SSB/k-1 =    105.9924
  
mean square within groups , MSW = SSW/N-k =    130.8520
  
F-stat = MSB/MSW =    0.8100
P value =   0.4634
  

-------------------

Ho: µ1=µ2=µ3
H1: not all means are equal
---------

SSTOT = 2174.765
SSBET = 211.985
SSW = 1962.780
dfBET = 2
dfW = 15
MSBET = 105.992
MSW = 130.852

value of the sample F statistic=0.810

the degrees of freedom?
(numerator)=2
(denominator)=15

ANOVA
Source of Variation SS df MS F P-value test decision
Between Groups 211.985 2 105.992 0.810 0.463 do not reject Ho
Within Groups 1962.780 15 130.852
Total 2174.765 17